# Versatile Motion-Language Models for Multi-turn Interactive Agents

### v2

@@ A.8 Detailed Explanation about Two-stage Baselines
TM2T 0.276 0.437 0.534 0.300 0.676 1.130
InterGEN 0.403 0.557 0.645 0.078 0.957 1.115
Next, we trained the text-to-motion diffusion model, InterGEN for the second stage. The performance of this model is reported in Table 9. InterGEN exhibits strong performance across all evaluation metrics, validating its effectiveness as the second stage in our two-stage approach. Specifically, InterGEN achieves Retrieval Precision scores of 0.403 (Top1), 0.557 (Top2), and 0.645 (Top3), which are substantially higher than those of the baseline MotionGPT (0.180, 0.262, 0.328) and our unified VIM model (0.318, 0.469, 0.568). Additionally, InterGEN excels in Diversity with a score of 0.957 and maintains a low Maximum Mean Discrepancy (MMDist) of 1.115, indicating high-quality and varied motion generation. Its FID score of 0.078 is notably competitive, reflecting the realism and coherence of the generated motions. These results validate the use of InterGEN as the second stage in our framework.
@@
In the motion reasoning task, the two-stage model integrates TM2T with large language models such as GPT-4o OpenAI (2024) and LLaMA-3.1-8B Dubey et al. (2024). Here, the motion components in the conversational data are replaced with captions generated by TM2T, which are then fed into the LLM for reasoning and response generation. In particular, the original input for the motion-language model was “USER:{question_1}, {motion_1}. ASSISTANT: {answer_1}. USER: {question_2}, {motion_2}.”, where the model has to predict “ASSISTANT: {answer_2}”. We replaced the motion into motion caption obtained by motion captioner for the input for LLM like “USER:{question_1}, {motion_caption_1}. ASSISTANT: {answer_1}. USER: {question_2}, {motion_caption_2}.”. Again, we utilized TM2T for the motion captioner mentioned in the previous section.
A.9 Template Forms for Pre-training and Instruction Tuning
Table 10: Template for Pretraining
@@ A.9 Template Forms for Pre-training and Instruction Tuning
Table 11: Template for Instruction Tunning
@@
We will detail the template forms utilized during the pre-training and instruction-tuning stages of our model development. Tables 10 and 11 illustrate the specific formats employed in each stage, providing a structured approach to aligning motion data with textual descriptions and enhancing the model’s interactive capabilities. All the templates are from MotionGPT (Jiang et al., 2023).
Pre-training Templates
@@
During the pre-training stage, our objective is to align motion and language representations by leveraging large language models (LLMs). We design tasks such as Text-to-Motion, Motion-to-Text, Reaction Generation, and Motion Prediction using paired datasets like InterX Xu et al. (2024a) and Interhuman Liang et al. (2024). The pre-training templates involve generating captions from motion sequences, creating motions based on textual descriptions, producing reaction motions in response to initial motions, and predicting subsequent motions from partial sequences, as summarized in Table 10. For single-person motion, we utilized text-to-motion, motion-to-text and motion prediction task during training.
Instruction-Tuning Templates
@@
In the instruction-tuning stage, we enhance the model’s ability to follow diverse instructions presented in a conversational format. Utilizing the INTER2-MT dataset alongside single-turn data from previous interactive motion datasets, we format user instructions and assistant responses to facilitate multi-turn interactions. Table 11 outlines the templates used for tasks such as generating motions from user prompts, describing motions based on user queries, and predicting motion continuations. By structuring the interactions in this manner, the model becomes adept at understanding and responding to various motion-related commands, thereby improving its performance in interactive scenarios.
A.10 Implementation Details
We set the codebook of the motion tokenizer as $K\in R^{512\times 512}$ for most comparisons, with residual depth $4$. The motion encoder $\mathcal{E}$ incorporates a temporal downsampling rate $l$ of 4. We utilize LLaMA-3.1-8B Dubey et al. (2024) as the underlying architecture for our language model. During the pertaining, we train the large language model (LLM) using a low-rank adaptor (LoRA) (Hu et al., 2022), including the embedding layer and the decoder head. The rank was set as $r=8$, $\alpha=16$, with the dropout rate set as $0.05$. During the instruction fine-tuning stage, we trained all the parameters. The learning rate was set as $0.0001$, and the warm-up ratio as $0.01$, the learning rate scheduler with cosine decay, and the AdamW optimizer.
A.11 More details about Evaluation Metric for Traditional Motion Related Tasks
Motion Quality
Motion Diversity
We have utilized diversity to evaluate the diversity of the motion following previous work (Jiang et al., 2023; Petrovich et al., 2023). To evaluate Diversity, the generated motions are split into two equal-sized subsets, and the Diversity metric is calculated as the average distance between motions within these subsets.
Condition Matching
TMR (Petrovich et al., 2023) offers motion/text feature extractors that produce geometrically coherent features for aligned text-motion pairs and vice versa. In this feature space, we evaluate motion-retrieval precision (R Precision) by combining the generated motion with 31 mismatched motions and calculating the top-1/2/3 matching accuracy between the text and motion. Furthermore, we assess the Multi-modal Distance (MM Dist), which measures the distance between the generated motions and their corresponding text.
A.12 User Subject Studies Protocols for motion editing
@@
We conducted user subject studies using the platform on the Mechanical Turk service from AWS (AWS, ).
A.12.1 Instructions
@@
The examples of ratings given to the user are shown in Figure 13.
Figure 13: The examples of ratings given to the user
A.12.2 Qualifying test
Before participating in the main user studies, all participants must pass a qualifying test to ensure they understand the evaluation criteria. In this test, participants are asked to assess four samples based on three metrics: Content Alignment, Fidelity of Motion, and Quality of Motion. Among the four samples, two are high-quality and derived from the ground-truth dataset, while the other two are low-quality—one is a mismatched motion with a single instruction, and the other is generated by the least effective model, MotionGPT. Participants must rate the low-quality samples lower than the high-quality ones in each of the three metrics. If any of the low-quality samples receive ratings that are equal to or higher than the high-quality samples in Content Alignment, Fidelity, or Quality of Motion, the participant will receive an error message and will need to adjust their ratings accordingly. This ensures that only participants who can accurately distinguish between high and low-quality motions based on the defined metrics proceed to the main study. The example of the qualifying test is demonstrated in Figure 14
Figure 14: Qualifying test in user subject studies
A.12.3 Detailed Survey Format
Figure 15: Caption
@@ A.12.3 Detailed Survey Format
A.13 Prompts for Data collection in Inter-$\text{MT}^{2}$
@@
We have utilized two different prompts in the data collection pipeline. One is generating two different motion captions with conversational data. The other one is generating one motion and conversational data based on the sample motion and corresponding caption from the base dataset, Inter-X (Xu et al., 2024a) and InterHuman Liang et al. (2024).
@@
Motion editing prompts without base sample is constructed as follows:
@@
You are an AI visual assistant, and you are seeing a motion. Design a conversation between you and a person building a conversation about editing this motion. In conversations, you should indicate who said using ”User:”, and”AI:” in the beginning but these two words do not occur in sentences. The answers should be in a tone that an AI visual assistant is seeing the motion and answering the question. The scenario should always contain two people in the scene. Generate a conversation about building a story from two different motions. The flow of the conversation is as follows: 1. Creating a scenario. REMBER to make a story in this. 2. Change the emotion or persona of just one person. 3. Describe how the motion will be changed, with one person maintaining the same motion. ”””Example: User: Let’s create a story starting from [Two individuals sitting across from each other, with one person extending his/her left hand and the other person extending their left hand. They proceed to participate in a wrist-wrestling competition]. AI: Two people are doing an arm-wrestling match, and each person is grabbing the right hand of the other person while sitting. User: The next scene is [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending both hands. They proceed to participate in a wrist-wrestling competition, where the second person utilizes both hands in an attempt to defeat the first person’s left hand.]. AI: The one person kept losing the game, which made him competitive to win the game.”””, ”””Example: User: Two friends are doing an arm-wrestling match. AI: [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending left hand. They proceed to participate in a wrist-wrestling competition] User: One person got competitive. AI: [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending both hands. They proceed to participate in a wrist-wrestling competition, where the second person utilizes both hands in an attempt to defeat the first person’s left hand.]. User: Explain the reason for the motion. AI: The one person kept losing the game, which made him cheat to win the game.”””, ”””Example: User: Two friends are doing an arm-wrestling match, like [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending left hand. They proceed to participate in a wrist-wrestling competition]. AI: Two people are doing an arm-wrestling match, each person is grabbing the right hand of the other person, while sitting. User: The one person kept losing the game, which made him competitive to win the game. Can you generate a motion of what would happen then? AI: [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending both hands. They proceed to participate in a wrist-wrestling competition, where the second person utilizes both hands in an attempt to defeat the first person’s left hand.]”””, ”””Example: User: Let’s start making a story. Two friends are doing an arm-wrestling match, like [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending their left hand. They proceed to participate in a wrist-wrestling competition]. AI: The one person kept losing the game, which made him competitive to win the game. User: Sounds interesting. Can you visualize it? AI: [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending both hands. They proceed to participate in a wrist-wrestling competition, where the second person utilizes both hands in an attempt to defeat the first person’s left hand.]””” =========== Example format for the [motion caption]: - One person approaches, raises his/her right hand to grab the other person’s right forearm, places his/her left hand on it, and walks in the direction the grabbed person is facing. - Two people face each other, one person lifts his/her right leg and walks towards the other person, stopping half a meter away. - A person falls and braces himself/herself on the ground with his/her right hand. Another person approaches, squats down, and grabs his/her left arm with both hands to assist him/her in standing up. The content inside the bracket ([]) is a caption for the motion. This is for visualizing the motion, which is not given in textual form during inference. I will denote this as [motion caption]. Please denote [motion caption] when AI or the user has to answer in the motion sequence.
@@
.. (continuing) Please make [motion caption] that is similar to the following action labels: [Action LABELS], and other motions like everyday routines (e.g., passing objects, greeting, communicating, etc.), and professional motions (e.g., Taekwondo, Latin dance, boxing, etc.) but still not necessary. Be creative too! Do not put [motion caption] in the same round, the user can also give motion to AI to reason from it too. Also, do not directly put [motion caption] twice in the round. You should put in only once, regarding both User and AI. [motion caption] are motion strings with skeleton information, which is for generating the motion. Do not repeat the caption. If you want to refer to these motions, just refer to it as the ’first motion’. But this motion string should be contained in the former to refer to. Try to make [motion caption] in details that do not require the previous context to generate the motion physically. ** Instead of the user fully describing what to do next, be more implicit, especially for the second motion, focusing more on the story. ** questions-answers not limited to the above examples. Questions should not be yes-no questions but wh-questions. The User-AI round should design at most 2. [motion caption] should appear only twice. Do not generate any new objects. Please follow the template from the example. It is better to keep the questions and answers concise. Try to be rational and keep in mind to make everything in sense, and the story smooth enough. Do not mention facial expressions or hands. Make the [motion caption] only ”twice” in the conversation. [motion caption] should always contain a description of two people. [motion caption] should have enough details for the motion, letting the model generate a correct motion by only accessing this caption without the previous context. Do not change the style of the motion caption. Do not make big and sudden changes in scenarios. REMEMBER: Try to make a description of the second motion that can be inferred by seeing the first motion. DO NOT GENERATE conversations that can be understandable without the previous context. FOCUS on **editing** the motion based on the emotion or personas. Users should NEVER ask AI to generate the motion giving details about what to do. LET AI infer about what to do based on the change of emotion. t is better to keep the questions and answers concise, with strictly following the format. Do not explain too much when generation motion. You are making a conversation about how the motion of the one person will change based on the persona, instead of keeping the story going on. The motion should be changed via body movement, not with facial expressions or hands. Do not directly [motion caption], this is just the format to guide you to fill the description there. Strictly follow the format. Generating **two** captions, with the changing persona for the motion. For the second caption, just change the motion of the second person. Do NOT LEAVE the [motion caption] holder! Do not put something like slightly, small, etc. It won’t be able to be visualized! Try to make a [motion caption] with the change of meaning of the motion, while maintaining a high-level scenario. Try to change the motion of the person dramatically, instead of changing just a few words.

### v5

@@ F.2 Text to Motion
task-specific approach
TM2T 0.276 0.437 0.534 0.300 0.676 1.130
MoMask 0.402 0.535 0.612 0.066 0.973 1.128
ComMDM 0.090 0.122 0.201 0.302 0.578 1.201
InterGEN 0.403 0.557 0.645 0.078 0.957 1.115
@@
[…]
F.3 Reaction Generation
Table F.5: Reaction Geneneration
Methods MPJPE $\downarrow$ FID $\downarrow$ Ret. Precision MMDist $\downarrow$
R Top1 $\uparrow$ R Top2 $\uparrow$ R Top3 $\uparrow$
@@ F.3 Reaction Generation
MoMask 1.602 0.112 0.109 0.328 0.412 1.178
unified approach
MotionGPT 3.441 0.355 0.079 0.104 0.355 1.246
[…]
VIM-w/o Inter-$\text{MT}^{2}$ 0.984 0.031 0.311 0.459 0.554 1.121
VIM (Ours) 0.690 0.019 0.381 0.537 0.625 1.110
Table F.5 presents a comparative analysis of various methods on the reaction generation task, focusing on metrics such as MPJPE, FID, Retrieval Precision, and MMDist. Our proposed unified approach, VIM, achieves the best performance across all evaluated metrics. Specifically, VIM attains the lowest MPJPE of 0.690 and the lowest FID of 0.019, indicating highly accurate joint position predictions and high fidelity in generated motions, respectively. In terms of retrieval precision, VIM outperforms both the task-specific method MoMask and other unified approaches, achieving R Top1, R Top2, and R Top3 scores of 0.381, 0.537, and 0.625. Additionally, VIM has the lowest MMDist of 1.110, suggesting that it generates motions closest to the real data distribution. The ablation model, VIM-w/o Inter-$\text{MT}^{2}$, also performs well but slightly lags behind VIM, highlighting the significance of the Inter-$\text{MT}^{2}$ component in enhancing performance. These results demonstrate that VIM not only narrows but effectively surpasses the performance gap between task-specific and unified approaches in reaction generation tasks, showcasing its effectiveness in generating accurate and realistic motion reactions.
G Limitations and Impact Statement
The expressiveness of our models remains limited when handling complex or previously unseen actions, indicating a need for further diverse motion source data in its ability to generalize across diverse motion scenarios. In addition, the sequence length becomes excessively long as we flatten the residual motion tokens, which can impact efficiency and computational resources. Leveraging additional transformer models to predict the residual token can reduce this work. Lastly, our method faces challenges in personalization and interpretability, as motion is inherently ambiguous and users may interpret the same motion in different ways. Addressing this issue will require incorporating more tailored approaches that adapt to individual user preferences and expectations through further human-in-the-loop feedback and refinement processes.
Broader Impact Statement
Our method opens up new possibilities for interactive motion modeling and understanding, potentially benefiting fields like robotics, virtual environments, and human-computer interaction. However, as the model evolves, careful consideration of ethical concerns, such as misinterpretation of motions or unintended behavioral biases, is crucial.
H Implementation Details
Figure H.4: Method Overview. Stage 1 involves training a motion tokenizer that encodes and decodes interactive motion data. In Stage 2, we pre-train the model by integrating motion and text data, allowing it to learn the alignment between text and motion. Stage 3 focuses on Instruction Tuning, fine-tuning the model to follow instructions and improve its responsiveness to conversational cues.
We set the codebook of the motion tokenizer as $K\in R^{512\times 512}$ for most comparisons, with residual depth $4$. The motion encoder $\mathcal{E}$ incorporates a temporal downsampling rate $l$ of 4. We utilize LLaMA-3.1-8B [9] as the underlying architecture for our language model. During the pertaining, we train the large language model (LLM) using a low-rank adaptor (LoRA) [17], including the embedding layer and the decoder head. The rank was set as $r=8$, $\alpha=16$, with the dropout rate set as $0.05$. During the instruction fine-tuning stage, we trained all the parameters. The learning rate was set as $0.0001$, the warm-up ratio as $0.01$, the learning rate scheduler with cosine decay, and the AdamW optimizer. For single-person modeling, we feed the encoder two identical copies of the motion. In other words, we concatenate the motion with itself so that the encoder, which expects two inputs, processes the same motion twice.
In this implementation, we first construct a motion tokenizer by training a motion encoder–decoder pair (Stage 1), where the encoder converts raw motion data into discrete tokens and the decoder can accurately reconstruct these motions from the tokens. Next, during pretraining (Stage 2), we fuse text tokens (from a large language model) with the motion tokens to jointly model their relationship, ensuring that the textual representations align with the motion embeddings. Finally, in the instruction-tuning phase (Stage 3), we refine the entire model using interactive text–motion tasks. Here, the model is trained to follow higher-level instructions by leveraging both the textual context and the learned motion tokens, enabling it to generate context-appropriate motions in response to user queries or commands. Figure H.4 illustrates the overall pipeline of the proposed method.
I Implementation details for Baselines
For training MotionGPT [18], TM2T [14] and MoMASK [15] in the interactive motion dataset, we have utilized the Flan-T5-base model [8] as a base large language model. We trained the model with Interhuman [22] and InterX [39] dataset, with the non-canonical representation, the same as the proposed method. To model the interactive motion, motion tokens of a person “A” and “B” are fed interleaved such as,
[…]
, where $k^{i;a}$ is the $i$-th token of motion $a$. Although scaling up the model can improve the performance, we conducted the experiment with the same base model as the original paper from MotionGPT [18] and Motionchain [19]. The original paper reported that increasing the model size did not significantly improve the model’s performance. We followed same motion token representation as MotionGPT for both TM2T [14] and MoMASK [15].
J Detailed Task Explanations
Motion Editing
@@ I Implementation details for Baselines
K Detailed Explanation about Two-stage Baselines
In Section 5.2 and Section 5.3, we have compared the proposed method with two-stage models. In particular, we have utilized TM2T [14] for the motion captioner and InterGEN [22] for the text-to-motion generation model.
K.1 Motion Editing
In the motion editing task, the two-stage approach first uses the motion-to-text (TM2T; [14]) model to generate a caption from the source motion and append the editing command. Then, the text-to-motion (InterGen; [22]) model produces the edited motion based on this caption and command. In particular, the input for the text-to-motion model is ”[motion caption]. [editing command]”.
We first trained the TM2T model with the InterHuman dataset [22]and the InterX [39] dataset, similarly to MotionGPT. Next, we trained the text-to-motion diffusion model, InterGEN for the second stage.
K.2 Motion Reasoning
In the motion reasoning task, the two-stage model integrates TM2T with large language models such as GPT-4o [27] and LLaMA-3.1-8B [9]. Here, the motion components in the conversational data are replaced with captions generated by TM2T, which are then fed into the LLM for reasoning and response generation. In particular, the original input for the motion-language model was “USER:{question_1}, {motion_1}. ASSISTANT: {answer_1}. USER: {question_2}, {motion_2}.”, where the model has to predict “ASSISTANT: {answer_2}”. We replaced the motion into motion caption obtained by motion captioner for the input for LLM like “USER:{question_1}, {motion_caption_1}. ASSISTANT: {answer_1}. USER: {question_2}, {motion_caption_2}.”. Again, we utilized TM2T for the motion captioner mentioned in the previous section.
L More details about Evaluation Metric for Traditional Motion Related Tasks
Motion Quality
Motion Diversity
We have utilized diversity to evaluate the diversity of the motion following previous work [18, 29]. To evaluate Diversity, the generated motions are split into two equal-sized subsets, and the Diversity metric is calculated as the average distance between motions within these subsets.
Text-Motion Matching
TMR [29] offers motion/text feature extractors that produce geometrically coherent features for aligned text-motion pairs and vice versa. In this feature space, we evaluate motion-retrieval precision (R Precision) by combining the generated motion with 31 mismatched motions and calculating the top-1/2/3 matching accuracy between the text and motion. Furthermore, we assess the Multi-modal Distance (MM Dist), which measures the distance between the generated motions and their corresponding text.
M Template Forms for Pre-training and Instruction Tuning
Table M.6: Template for Pretraining
@@
Table M.7: Template for Instruction Tunning
@@
We will detail the template forms utilized during the pre-training and instruction-tuning stages of our model development. Tables M.6 and M.7 illustrate the specific formats employed in each stage, providing a structured approach to aligning motion data with textual descriptions and enhancing the model’s interactive capabilities. All the templates are from MotionGPT [18].
M.1 Pre-training Templates
@@
During the pre-training stage, our objective is to align motion and language representations by leveraging large language models (LLMs). We design tasks such as Text-to-Motion, Motion-to-Text, Reaction Generation, and Motion Prediction using paired datasets like InterX [39] and Interhuman [22]. The pre-training templates involve generating captions from motion sequences, creating motions based on textual descriptions, producing reaction motions in response to initial motions, and predicting subsequent motions from partial sequences, as summarized in Table M.6. For single-person motion, we utilized text-to-motion, motion-to-text and motion prediction task during training.
M.2 Instruction-Tuning Templates
@@
In the instruction-tuning stage, we enhance the model’s ability to follow diverse instructions presented in a conversational format. Utilizing the INTER2-MT dataset alongside single-turn data from previous interactive motion datasets, we format user instructions and assistant responses to facilitate multi-turn interactions. Table M.7 outlines the templates used for tasks such as generating motions from user prompts, describing motions based on user queries, and predicting motion continuations. By structuring the interactions in this manner, the model becomes adept at understanding and responding to various motion-related commands, thereby improving its performance in interactive scenarios.
N Data Sample Visualization
The samples from the synthesized dataset, Inter-$\text{MT}^{2}$, are illustrated in Figure N.5.
Figure N.5: Sample from Inter-$\text{MT}^{2}$ dataset. The left column visualizes samples of motion editing, and the right column shows examples from the motion reasoning task.
O Inter-$\text{MT}^{2}$ Statistics
Table O.8: Statistics on Inter-$\text{MT}^{2}$.
Total Train Val. Test
# of Samples 82736 66194 4141 12401
# of Motions 317749 132388 8282 24802
From Dataset 56395 50258 3142 2995
Synthesized 96676 82130 5140 9406
We collected 82K samples of multi-turn conversational data, each involving interactive motions. Of these, 30K samples focus on motion editing, 30K on reasoning about past or future scenarios, and 12K on story generation. Each sample includes four to eight conversation turns and two distinct motions. The dataset contains 96K motions generated using a text-to-motion diffusion model, while 56K motions come from the original source dataset. The train-validation-test set is randomly split by the ratio 0.8:0.05:0.15.
P Qualitative Results
We visualize our result gallery on motion editing in Figure P.6 and on motion reasoning in Figure P.7. Furthermore, the results for motion-to-text (Figure P.8), text-to-motion (Figure P.9), and reaction generation (Figure P.10) are demonstrated. In figure P.11, we demonstrated the generation ability of the proposed method in longer contexts with a failure case.
Figure P.6: Generated motion from source and editing command.
Figure P.7: Generated responses based on the previous conversations for motion reasoning task.
Figure P.8: Motion-to-text results. The blue part is generated motion captions from source motions.
Figure P.9: Text-to-motion results. The blue part is generated motions from the motion caption.
Figure P.10: Reaction Generation. The input motion is orange, while the generated reactive motion is colored blue.
Figure P.11: We demonstrate the ability of the proposed method to generate responses in long conversations and also the failure cases.
Q User Subject Studies Protocols for motion editing
@@
We conducted user subject studies using the platform on the Mechanical Turk service from AWS [3].
Q.1 Instructions
@@
The examples of ratings given to the user are shown in Figure Q.12.
Figure Q.12: The examples of ratings given to the user
Q.2 Qualifying test
Before participating in the main user studies, all participants must pass a qualifying test to ensure they understand the evaluation criteria. In this test, participants are asked to assess four samples based on three metrics: Content Alignment, Fidelity of Motion, and Quality of Motion. Among the four samples, two are high-quality and derived from the ground-truth dataset, while the other two are low-quality—one is a mismatched motion with a single instruction, and the other is generated by the least effective model, MotionGPT. Participants must rate the low-quality samples lower than the high-quality ones in each of the three metrics. If any of the low-quality samples receive ratings that are equal to or higher than the high-quality samples in Content Alignment, Fidelity, or Quality of Motion, the participant will receive an error message and will need to adjust their ratings accordingly. This ensures that only participants who can accurately distinguish between high and low-quality motions based on the defined metrics proceed to the main study. The example of the qualifying test is demonstrated in Figure Q.13
Figure Q.13: Qualifying test in user subject studies
Q.3 Detailed Survey Format
Figure Q.14: Survey Example
@@
R Prompts for Data collection in Inter-$\text{MT}^{2}$
@@
We have utilized two different prompts in the data collection pipeline. One is generating two different motion captions with conversational data. The other one is generating one motion and conversational data based on the sample motion and corresponding caption from the base dataset, Inter-X [39] and InterHuman [22].
@@
Motion editing prompts without base samples is constructed as follows:
@@
You are an AI visual assistant, and you are seeing a motion. Design a conversation between you and a person building a conversation about editing this motion. In conversations, you should indicate who said using ”User:”, and”AI:” in the beginning but these two words do not occur in sentences. The answers should be in a tone that an AI visual assistant is seeing the motion and answering the question. The scenario should always contain two people in the scene. Generate a conversation about building a story from two different motions. The flow of the conversation is as follows: 1. Creating a scenario. REMBER to make a story in this. 2. Change the emotion or persona of just one person. 3. Describe how the motion will be changed, with one person maintaining the same motion. ”””Example: User: Let’s create a story starting from [Two individuals sitting across from each other, with one person extending his/her left hand and the other person extending their left hand. They proceed to participate in a wrist-wrestling competition]. AI: Two people are doing an arm-wrestling match, and each person is grabbing the right hand of the other person while sitting. User: The next scene is [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending both hands. They proceed to participate in a wrist-wrestling competition, where the second person utilizes both hands in an attempt to defeat the first person’s left hand.]. AI: The one person kept losing the game, which made him competitive to win the game.”””, ”””Example: User: Two friends are doing an arm-wrestling match. AI: [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending left hand. They proceed to participate in a wrist-wrestling competition] User: One person got competitive. AI: [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending both hands. They proceed to participate in a wrist-wrestling competition, where the second person utilizes both hands in an attempt to defeat the first person’s left hand.]. User: Explain the reason for the motion. AI: The one person kept losing the game, which made him cheat to win the game.”””, ”””Example: User: Two friends are doing an arm-wrestling match, like [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending left hand. They proceed to participate in a wrist-wrestling competition]. AI: Two people are doing an arm-wrestling match, each person is grabbing the right hand of the other person, while sitting. User: The one person kept losing the game, which made him competitive to win the game. Can you generate a motion of what would happen then? AI: [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending both hands. They proceed to participate in a wrist-wrestling competition, where the second person utilizes both hands in an attempt to defeat the first person’s left hand.]”””, ”””Example: User: Let’s start making a story. Two friends are doing an arm-wrestling match, like [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending their left hand. They proceed to participate in a wrist-wrestling competition]. AI: The one person kept losing the game, which made him competitive to win the game. User: Sounds interesting. Can you visualize it? AI: [Two individuals sit across from each other, with one person extending his/her left hand and the other person extending both hands. They proceed to participate in a wrist-wrestling competition, where the second person utilizes both hands in an attempt to defeat the first person’s left hand.]””” =========== Example format for the [motion caption]: - One person approaches, raises his/her right hand to grab the other person’s right forearm, places his/her left hand on it, and walks in the direction the grabbed person is facing. - Two people face each other, one person lifts his/her right leg and walks towards the other person, stopping half a meter away. - A person falls and braces himself/herself on the ground with his/her right hand. Another person approaches, squats down, and grabs his/her left arm with both hands to assist him/her in standing up. The content inside the bracket ([]) is a caption for the motion. This is for visualizing the motion, which is not given in textual form during inference. I will denote this as [motion caption]. Please denote [motion caption] when AI or the user has to answer in the motion sequence. Please make [motion caption] that is similar to the following action labels: [Action LABELS], and other motions like everyday routines (e.g., passing objects, greeting, communicating, etc.), and professional motions (e.g., Taekwondo, Latin dance, boxing, etc.) but still not necessary. Be creative too! Do not put [motion caption] in the same round, the user can also give motion to AI to reason from it too. Also, do not directly put [motion caption] twice in the round. You should put in only once, regarding both User and AI. [motion caption] are motion strings with skeleton information, which are used to generate motion. Do not repeat the caption. If you want to refer to these motions, just refer to it as the ’first motion’. But this motion string should be contained in the former to refer to. Try to make [motion caption] in details that do not require the previous context to generate the motion physically. ** Instead of the user fully describing what to do next, be more implicit, especially for the second motion, focusing more on the story. ** questions-answers not limited to the above examples. Questions should not be yes-no questions but wh-questions. The User-AI round should design at most 2. [motion caption] should appear only twice. Do not generate any new objects. Please follow the template from the example.
@@
It is better to keep the questions and answers concise. Try to be rational and keep in mind to make everything in sense, and the story smooth enough. Do not mention facial expressions or hands. Make the [motion caption] only ”twice” in the conversation. [motion caption] should always contain a description of two people. [motion caption] should have enough details for the motion, letting the model generate a correct motion by only accessing this caption without the previous context. Do not change the style of the motion caption. Do not make big and sudden changes in scenarios. REMEMBER: Try to make a description of the second motion that can be inferred by seeing the first motion. DO NOT GENERATE conversations that can be understandable without the previous context. FOCUS on **editing** the motion based on the emotion or personas. Users should NEVER ask AI to generate the motion giving details about what to do. LET AI infer about what to do based on the change of emotion. t is better to keep the questions and answers concise, with strictly following the format. Do not explain too much when generation motion. You are making a conversation about how the motion of the one person will change based on the persona, instead of keeping the story going on. The motion should be changed via body movement, not with facial expressions or hands. Do not directly [motion caption], this is just the format to guide you to fill the description there. Strictly follow the format. Generating **two** captions, with the changing persona for the motion. For the second caption, just change the motion of the second person. Do NOT LEAVE the [motion caption] holder! Do not put something like slightly, small, etc. It won’t be able to be visualized! Try to make a [motion caption] with the change of meaning of the motion, while maintaining a high-level scenario. Try to change the motion of the person dramatically, instead of changing just a few words.